identify the vertex, focus, and directrix of the problem. y=1/2(x-1)^2
Answer by stanbon(75887)

identify the vertex, focus, and directrix of the problem. y=1/2(x-1)^2
----
Form: (x-h)^2 = 4p(y-k)
----
Your problem:
(x-1)^2 = 2(y-0)
-----
Vertex: (1,0)
----
4p = 2
p = 1/2
---------
Focus: (1, 0+1/2) = (1,1/2)
Directix: y = -1/2
======================
Cheers,
Stan H.